OpenStudy (anonymous):

Simplify square root of 2 multiplied by the cube root of 2.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sqrt{2}\times \sqrt[3]{2} }\) Like this?

OpenStudy (anonymous):

yeap

OpenStudy (solomonzelman):

There is are some rules you need to apply. \(\color{#000000 }{ \displaystyle \sqrt[\Large \color{blue}{c}]{a^\color{red}{b}} = a^{\color{red}{b}/\color{blue}{c}} }\) \(\color{#000000 }{ \displaystyle a^\color{blue}{c}\times a^\color{red}{b}=a^{\color{blue}{c}+\color{red}{b}} }\)

OpenStudy (solomonzelman):

I am lagging a little-:(

OpenStudy (anonymous):

same

OpenStudy (anonymous):

so 2\[2^{1}\]x\[2^{3}\]=\[2^{1+3}\]?

OpenStudy (solomonzelman):

Hint: \(\color{#000000 }{ \displaystyle \sqrt{2}=\sqrt[2]{2} }\) (the little number befor the root, above, is always 2 when not specified) \(\color{#000000 }{ \displaystyle \sqrt{2}=\sqrt[2]{2}=\sqrt[2]{2^1} }\) (when the number is multiplied one time times itself, that is - when it is written only once, that is same as rasing to the power of a 1) And by this principle, \(\color{#000000 }{ \displaystyle \sqrt[3]{2}=\sqrt[3]{2^1} }\)

OpenStudy (anonymous):

ohh ok